# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ 4.1 $\pi^{\pm}$ condensate nuclear systems
In Fig. 5 by solid curves we show the binding energy per nucleon for the purely neutron matter (PNM). Other curves show calculations for the isospin symmetric matter (ISM) with and without the developed condensate. On left panel the nucleon part (dashed line) is calculated with a soft Friedman-Pandharipande (FP) equation of state Friedman and Pandharipande (1981) and on right panel with a stiffer RMF equation of state, see Eq. (15). In the RMF case parameters are tuned to describe the saturation density $n_{0}=0.160$ fm-3, the binding energy $\mathcal{E}_{0}=-16.0$ MeV, the compressibility modulus $K=273.6$ MeV, the effective nucleon mass $m_{N}^{*}(n_{0})=0.755$ MeV, and the value of the coefficient of the nuclear symmetry energy at saturation, $J=32.0$ MeV. With these parameters the maximum neutron star mass is $2.09\,M_{\odot}$, being larger $2\,M_{\odot}$ in agreement with the modern experimental data. The attraction strength of the $\pi^{\pm}$ condensate contribution, $E_{\pi}$ in Eq. (28), is controlled by the value of the $g^{\prime}$ Landau-Migdal parameter. The density dependence of this parameter is poorly known, therefore following Voskresensky et al. (1977); Migdal et al. (1990) we allow for its variation. So, the energies per nucleon of the baryon, $pn\Delta$, matter with the pion condensate (dash-dotted lines) are shown for several values of $g^{\prime}$. Note that various fits of the atomic nuclei data produce the value of $g^{\prime}(n_{0})\approx 0.7$. At this instance to avoid a possible misunderstanding we should mention that different authors use different normalizations of $g^{\prime}$ and distinct expressions for the pion polarization operator. The smaller $g^{\prime}(n)$ is the stronger is the pion condensate attraction for $n>n_{c}^{p}$. In Migdal et al. (1990) there were given some arguments for a possible decrease of $g^{\prime}(n)$ with increasing density. For the FP equation of state (left panel in Fig. 5) the energy of the pion condensate matter becomes negative in some interval of density for $g^{\prime}(n>n_{c}^{p})<0.482$. In this case there may exist anomalous superheavy pion condensate nuclei, nuclearites and nuclei-stars of arbitrary size bounded by nuclear forces rather than by gravity (for $10^{3}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}A\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{56}$), as it was suggested in Ref. Voskresensky et al. (1977). The maximal net charge of superheavy nucleus with $A\gg 1/e^{3}$ observed at infinity behaves as $A^{1/3}$. In this case the extra charge is with necessity screened by electrons and muons occupying the vacuum shell. The pion condensate state becomes an absolute ground state on nuclear matter if $g^{\prime}<0.446$, then $\min_{n}\{{E}_{B+\pi}\}<-16$ MeV. For $0.482<g^{\prime}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.6$ there is a minimum in energy ${E}_{B+\pi}$ that would correspond to a metastable state. For the RMF nucleon equation of state (right panel in Fig. 5) the anomalous states may exist for $g^{\prime}<0.426$ and the absolute ground sate of matter appears for $g^{\prime}<0.381$. The metastable states are possible for $0.426<g^{\prime}<0.492$.
4.2 Superheavy $\pi^{0}$-condensate ferromagnetic nuclear systems
@@
Possibility of neutron anomalous $\pi^{0}$ condensate nuclei was discussed in Migdal (1974); Migdal et al. (1976). Ferromagnetic behavior of $\pi^{0}$ condensate was studied within the $\sigma$ model first in Dautry and Nyman (1979). The $\sigma\pi^{0}$ running wave condensate in the neutron matter corresponds to a state, in which all neutron spins are aligned, presumably with some macroscopic domain structure. More recently there were found that extra contributions to the $p$ wave $\pi^{0}$ condensate term in the Lagrangian may arrive from the so-called Wess-Zumino-Witten axial anomaly term describing the anomalous interaction of the neutral pion field with the external electromagnetic field, and a related pion contribution to the baryon current, cf. Son and Stephanov (2008); Eto et al. (2013); Hashimoto (2015); Brauner and Yamamoto (2017); Huang et al. (2018).
Comparison of the ferromagnetic phase of the $\pi^{0}$ condensate appearing due to the axial anomaly term and the alternating layer structure of the pion condensate Takatsuka et al. (1978) demonstrated that the latter state is probably energetically more favorable Hashimoto (2015). These effects were recently studied in Ref. Voskresensky (2025).
4.3 About $s$ wave pion condensation
@@
Pion off-mass-shell effects are of primary importance for description of the pion spectra and a possibility of the pion condensation. The most important contribution to the pion polarization operator is given by the $p$ wave pion-nucleon and the pion-$\Delta$-isobar interactions. For the $s$ wave part of the pion polarization operator already first works employed the Weinberg-Tomozawa expression, $\Pi^{s}_{\pi^{-}}=C(n_{n}-n_{p})\omega$, with $C\simeq 1/m^{2}_{\pi}$, which does not contribute in case of the ISM. Parametrization of the optical pion-nucleus potential Migdal (1978); Migdal et al. (1990); Voskresensky (1993) used the fully off-mass-shell pion-nucleon amplitude, which fulfills the current algebra theorems and the canonical PCAC condition (see the MSTV model in Voskresensky (2025)). Oppositely, Refs. Delorme et al. (1992); Ericson (1994); Kolomeitsev et al. (2003a, b) (see the KKW model in Voskresensky (2025)) used the on-mass-shell pion-nucleon amplitude, taking incoming and outgoing pion 4-momenta such that $q^{2}=q^{{}^{\prime}2}=m^{2}_{\pi}$, that does not allow to fulfill the so called Adler and Weinberg current algebra conditions. Subtleties associated with the current algebra theorems and field redefinitions were discussed in Voskresensky (2022). The $s$ wave pion-nucleon interactions in the linear sigma model, and in the Manohar-Georgi and Gasser-Sainio-Svarc models with finite number of terms in Lagrangians, as well as in a general phenomenological approach were reviewed.
@@
With the MSTV model for the $s$ wave pion-nucleon interaction the $s$ wave pion condensation in the ISM does not occur at least up to very high densities. Oppositely, if the KKW model Delorme et al. (1992); Ericson (1994); Kolomeitsev et al. (2003a, b) were valid, the $s$ wave pion condensation in the ISM would be expected to occur already for $n>n_{c}^{s}\sim(1.4-2.5)n_{0}$, if the nucleon-nucleon correlations are not taken into account, resulting in the energy density gain
$E_{\pi}^{\rm KKW}(n)\simeq-\frac{m^{4}_{\pi}}{2\lambda}\left(\frac{n}{n_{c}^{s}}-1\right)^{2}\Theta(n-n_{c}^{s})\,.$ (30)
@@
In Fig. 6 we plot the energy per particle of the ISM with and without the contribution of the $s$ wave pion condensation (30). We use the same RMF model as on the right panel of Fig. 5. The pion-pion self interaction parameter is chosen to be $\lambda=m_{\pi}^{2}/(2f_{\pi}^{2})$. The value of the softening of the equation of state essentially depends on the value of the critical density. As we see, for $n_{c}^{s}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}2.5\,n_{0}$ the softening is weak, whereas for $n_{c}^{s}\leq 1.55\,n_{0}$ the effect is strong and there appears possibility of a metastable state with the $s$ wave pion condensate. Thereby one could expect to observe experimental manifestations of the $s$ wave pion condensate in heavy-ion collisions. We must, however, emphasize that the estimate of the value of the critical density $n_{c}^{s}$ remains strongly model-dependent even within the on-mass shell based model Delorme et al. (1992); Ericson (1994); Kolomeitsev et al. (2003a, b). For example, the $NN$ correlation effects, which cannot be constrained from the experiments at present, could shift the condensation critical point to higher values Voskresensky (2025).
@@
5 $\Delta$ resonance matter
The idea of a stable or metastable baryon-resonance matter was introduced in Ref. Troitsky and Chekunaev (1979). The Fermi momentum of the nucleon is $p_{\rm F}\propto n^{1/3}/\nu^{1/3}$, the total Fermi energy of baryons grows as $E_{\rm F}\propto n^{2/3}\nu^{2/3}$ where $\nu$ is the degeneracy factor. When the nucleon Fermi energy, $E_{\rm F}$, reaches the difference between effective masses $m^{*}_{\Delta}-m^{*}_{N}$, it becomes energetically favorable to create $\Delta$ isobars. The appearance of any new degree of freedom results in a softening of the equation of state. The nucleon–nucleon-hole contribution to the pion polarization operator is $\Pi\propto-\nu p_{\rm F}k^{2}\propto\nu^{2/3}n^{1/3}$, see Voskresensky and Mishustin (1982). Therefore, the latter quantity effectively corresponds to the $\nu^{2}$ times higher density compared to the case of $\nu=1$. The extra $\Delta$$\Delta$-hole attractive term to the pion polarization operator leads to a further decrease of the critical density of the pion condensation. Moreover, let us recall that the spin-isospin degeneracy factor for $\Delta$ is $\nu=16$ compared to 4 for nucleons and that $f_{\pi N\Delta}\simeq 2f_{\pi NN}$.
$\Delta$ isobars are produced copiously in relativistic heavy-ion collisions forming a resonance matter Metag (1993). A possibility of density isomer states was discussed in Boguta (1982); Waldhauser et al. (1987); Cubero et al. (1988); Diakonov and Mirlin (1988) within RMF models and the chiral model of the nucleon.
@@ 5 $\Delta$ resonance matter
6 Strangelets, hybrid and strange stars
6.1 Various models
Discussions of hypothetical strangelets and hybrid and quark stars began with the works Ivanenko and Kurdgelaidze (1965); Itoh (1970); Bodmer (1971); Witten (1984); Farhi and Jaffe (1984) and remain a hot topic till these days, see, e.g., Refs. Weber (2005); Farrell et al. (2024); Lugones and Grunfeld (2025); Di Clemente et al. (2025) and references therein. In Ref. Witten (1984) Witten estimated the energy per quark for a three-flavour system (strange quark matter) to be $\sim 90\%$ of the energy of a two-flavour system. This additional energy gain may compensate the energy loss due to the heavier strange quark mass leading, thereby, to stable quark matter.
The most commonly employed model of quark matter is the MIT bag model Chodos et al. (1974). By balancing the vacuum pressure outside the bag with the pressure of quarks inside it, one mimics phenomena of quark confinement and asymptotic freedom Farhi and Jaffe (1984). Besides the MIT-bag based approaches, the Nambu–Jona-Lasinio model is often used, see Ref. Klevansky (1992). The latter does not describe confinement of quarks and its preferred sets of parameters do not generally predict stable strange quark matter Buballa and Oertel (1999). The RMF approach proposed in Ref. Kaltenborn et al. (2017) gives a simple way to model confinement of quarks via introducing the particular density dependence of quark masses divergent for low baryon densities and a density-dependent screening effect. The screening is described analogously to an excluded volume effect for a hadronic matter model with density-dependent coupling constants. Effects of various pasta phases occurring in the quark-hadron phase transition Voskresensky et al. (2003) were then investigated in Ref. Maslov et al. (2019) for a set of RMF equations of state for both the hadron matter and the quark one. These approaches allow for hybrid stars but exclude stable quark star configurations. Some constraints may come from the lattice calculations as was discussed in Ivanov et al. (2005); Peshier et al. (2002); Schulze and Kampfer (2009). As a possible result, families of ordinary neutron stars and pure quark stars can be well separated. Within these models, the quark stars have low masses ($M\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1M_{\odot}$), high densities ($n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10n_{0}$), and small radii ($R\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}6$ km). Also, one uses the quark models with quark masses varying with the baryon number density, see Xia et al. (2014). Situation remains rather uncertain. Some works give arguments that even two-flavor quark matter can be stable or metastable, other argue against stable quark matter, e.g. see Holdom et al. (2018); Bai and Chen (2025) and references therein.
The quark objects are anticipated to be color superconductors. Various quark-paring patterns were investigated: color-flavor-locked phase, 2SC phase, spin-locking phase etc., see Alford et al. (2008); Farrell et al. (2024). Their effect on the equation of state is not significant even if the pairing gap is as large as $\Delta\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}100$ MeV. Owing a strong interaction a fluctuation region near critical point of the color superconductivity was estimated in Voskresensky (2004) as a broad and it is not excluded that some effects of the color superconducting fluctuations can be manifested in heavy-ion collisions.

### v2

@@ IV.2 $\pi^{\pm}$ condensate nuclear systems
In Fig. 7 by solid curves we show the binding energy per nucleon for the purely neutron matter (PNM). Other curves show calculations for the ISM with and without the developed condensate done with the help of Eq. (32). Actually this expression does not hold for $n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n_{c}^{(p)}$ since it was derived for $n\gg n_{c}^{(p)}$, see a discussion in [210]. However we conjecture that for $n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}(2-3)n_{0}$ we already can deal with the developed pion condensate On left panel the nucleon part (dashed line) is calculated with a soft Friedman-Pandharipande (FP) equation of state [92] and on right panel with a stiffer RMF equation of state, see Eq. (15). In the RMF case parameters are tuned to describe the saturation density $n_{0}=0.160$ fm-3, the binding energy $\mathcal{E}_{0}=-16.0$ MeV, the compressibility modulus $K=273.6$ MeV, the effective nucleon mass $m_{N}^{*}(n_{0})=0.755$ MeV, and the value of the coefficient of the nuclear symmetry energy at saturation, $J=32.0$ MeV. With these parameters the maximum neutron star mass is $2.09\,M_{\odot}$, being larger $2\,M_{\odot}$ in agreement with the modern experimental data. The attraction strength of the $\pi^{\pm}$ condensate contribution, $E_{\pi}$ in Eq. (32), is controlled by the value of the $g^{\prime}$ Landau-Migdal parameter. The density dependence of this parameter is poorly known, therefore following [280, 210] we allow for its variation. So, the energies per nucleon of the baryon, $pn\Delta e\mu$ matter with the pion condensate (dash-dotted lines) are shown for several values of $g^{\prime}$. Note that various fits of the atomic nuclei data produce the value of $g^{\prime}(n_{0})\approx 0.7$. At this instance to avoid a possible misunderstanding we should mention that different authors use different normalizations of $g^{\prime}$ and distinct expressions for the pion polarization operator. The smaller $g^{\prime}(n)$ is the stronger is the pion condensate attraction for $n>n_{c}^{\rm(p)}$. In [210] there were given some arguments for a possible decrease of $g^{\prime}(n)$ with increasing density. For the FP equation of state (left panel in Fig. 7) the energy of the pion condensate matter becomes negative in some interval of density for $g^{\prime}(n>n_{c}^{\rm(p)})<0.48$. In this case there may exist superdense abnormal pion condensate nuclei, nuclearites and nuclei-stars of arbitrary size bounded by nuclear forces rather than by gravity (for $10^{3}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}A\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}10^{56}$), as it was suggested in Ref. [280]. The maximal net charge of an abnormal nucleus with $A\gg 1/e^{3}$ observed at infinity behaves as $A^{1/3}$. In this case the extra charge is with necessity screened by electrons and muons occupying the vacuum shell. The pion condensate state becomes an absolute ground state of nuclear matter if $g^{\prime}<0.45$, then $\min_{n}\{{E}_{B+\pi}\}<-16$ MeV. For $0.48<g^{\prime}<0.62$ there is a minimum in energy ${E}_{B+\pi}$ that would correspond to a positive energy metastable state and for $0.45<g^{\prime}<0.48$ to a negative energy metastable state. For the RMF nucleon equation of state, which we used (see right panel in Fig. 7) the abnormal negative energy states may exist for $g^{\prime}<0.43$ and the absolute ground state of matter appears for $g^{\prime}<0.38$. The metastable states with positive energy are possible for $0.43<g^{\prime}<0.63$.
IV.3 Dense $\pi^{0}$-condensate ferromagnetic nuclear systems
@@
Possibility of neutron-enriched abnormal $\pi^{0}$ condensate nuclei was discussed in [214, 211]. Ferromagnetic behavior of $\pi^{0}$ condensate was studied within the $\sigma$ model first in [62]. The $\sigma\pi^{0}$ running wave condensate in the neutron matter corresponds to a state, in which all neutron spins are aligned, presumably with some macroscopic domain structure. Recently there were found that extra contributions to the p-wave $\pi^{0}$ condensate term in the Lagrangian may arrive from the so-called Wess-Zumino-Witten axial anomaly term describing the anomalous interaction of the neutral pion field with the external electromagnetic field and the rotation, and a related pion contribution to the baryon current, cf. [264, 78, 116, 40, 125]. Comparison of the ferromagnetic phase of the $\pi^{0}$ condensate appearing due to the axial anomaly term and the alternating layer structure of the pion condensate [265] demonstrated that the latter state is probably energetically more favorable [116]. These effects were recently studied in Ref. [290].
IV.4 About the s-wave pion condensation
@@
Pion off-mass-shell effects are of primary importance for description of the pion spectra and a possibility of the pion condensation. The most important contribution to the pion polarization operator is given by the p-wave pion-nucleon and the pion-$\Delta$-isobar interactions. For an s-wave part of the pion polarization operator already first works employed the Weinberg-Tomozawa expression, $\Pi^{s}_{\pi^{-}}=C(n_{n}-n_{p})\,\omega$, with $C\simeq 1/(2f_{\pi}^{2})\simeq 1/m^{2}_{\pi}$, which does not contribute in case of the ISM. Parametrization of the optical pion-nucleus potential [215, 210, 284] used the fully off-mass-shell pion-nucleon amplitude, which fulfills the current algebra theorems and the canonical PCAC condition (see the MSTV model in [290]). Oppositely, Refs. [66, 77, 162, 150] (see the KKW model in [290]) used the on-mass-shell pion-nucleon amplitude, taking incoming and outgoing pion 4-momenta such that $q^{2}=q^{{}^{\prime}2}=m^{2}_{\pi}$, that does not allow to fulfill the so called Adler and Weinberg current algebra conditions. Subtleties associated with the current algebra theorems and field redefinitions were discussed in [287]. The s-wave pion-nucleon interactions in the linear sigma model, and in the Manohar-Georgi and Gasser-Sainio-Svarc models with finite number of terms in Lagrangians, as well as in a general phenomenological approach were reviewed.
@@
With the MSTV model for the s-wave pion-nucleon interaction the s-wave pion condensation in the ISM does not occur at least up to very high densities. Oppositely, if the KKW model [66, 77, 162, 150] were valid, the s-wave pion condensation in the ISM would be expected to occur already for $n>n_{c}^{\rm(s)}\sim(1.4-2.5)n_{0}$, if the nucleon-nucleon correlations are not taken into account, resulting in the energy density gain
$E_{\pi}^{\rm KKW}(n)\simeq-\frac{m^{4}_{\pi}}{2\lambda}\left(\frac{n}{n_{c}^{\rm(s)}}-1\right)^{2}\Theta(n-n_{c}^{\rm(s)})\,.$ (34)
@@
In Fig. 8 we plot the energy per particle of the ISM with and without contribution of the s-wave pion condensation (34). We use the same RMF model as on the right panel of Fig. 7. The pion-pion self interaction parameter is chosen to be $\lambda=m_{\pi}^{2}/(2f_{\pi}^{2})$. The value of the softening of the equation of state essentially depends on the value of the critical density. As we see, for $n_{c}^{\rm(s)}\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}2.5\,n_{0}$ the softening is weak, whereas for $n_{c}^{\rm(s)}\leq 1.55\,n_{0}$ the effect is strong and there appears possibility of a metastable state with the s-wave pion condensate. Thereby one could expect to observe experimental manifestations of the s-wave pion condensate in heavy-ion collisions. We must, however, emphasize that the estimate of the value of the critical density $n_{c}^{\rm(s)}$ remains strongly model-dependent even within the on-mass shell based model [66, 77, 162, 150]. For example, the $NN$ correlation effects in the $s$-wave pion polarization operator, which cannot be constrained from the experiments at present, could shift the condensation critical point to higher values [290].
@@
V $\Delta$ resonance matter
Above we indicated that the pion-nucleon-$\Delta$ attraction promotes a decrease of the effective pion gap and occurrence of the pion condensation at smaller baryon density. Now we will review the idea of $\Delta$ resonance matter.
The idea of a stable or metastable baryon-resonance matter was introduced in Ref. [269]. The Fermi momentum of the nucleon is $p_{\rm F}\propto n^{1/3}/\nu^{1/3}$, the total Fermi energy of baryons grows as $E_{\rm F}\propto n^{2/3}\nu^{2/3}$ where $\nu$ is the degeneracy factor. When the nucleon Fermi energy, $E_{\rm F}$, reaches the difference between effective masses $m^{*}_{\Delta}-m^{*}_{N}$, it becomes energetically favorable to create $\Delta(1232)$ isobars. The appearance of any new degree of freedom results in a softening of the equation of state. The nucleon–nucleon-hole contribution to the pion polarization operator is $\Pi_{\pi NN}\propto-\nu p_{\rm F}k^{2}\propto\nu^{2/3}n^{1/3}$, Therefore, the latter quantity effectively corresponds to the $\nu^{2}$ times higher density compared to the case of $\nu=1$. The $\Delta$–nucleon-hole contribution is $\Pi_{\pi\Delta N}\propto-nk^{2}$, see [276]. The extra $\Delta$$\Delta$-hole attractive term to the pion polarization operator leads to a further decrease of the critical density of the pion condensation. Moreover, let us recall that the spin-isospin degeneracy factor for $\Delta$ is $\nu=16$ compared to 4 for nucleons and that $f_{\pi N\Delta}\simeq 2f_{\pi NN}$.
@@ V $\Delta$ resonance matter
$\Delta$ isobars are produced copiously in relativistic heavy-ion collisions forming a resonance matter [203]. A possibility of density isomer states was discussed in [36, 294, 59, 68] within RMF models and the chiral model of the nucleon.
@@
Note that the occupation of the new $\Delta$ Fermi sea occurs at nucleon densities exceeding a critical value $n>n_{c}\sim(2-4)n_{0}$, which depends on the $N/Z$ ratio and parameters of the model under consideration. Appearance of $\Delta$ isobars makes the nuclear matter more stable. Besides this, the occupation of the $\Delta$ Fermi sea additionally promotes a decrease of the effective pion gap and a pion condensation at smaller baryon density.
In addition to the occupation of the $\Delta$ Fermi sea, in the neutron star matter (at $N\gg Z$), $\Lambda$, $\Sigma^{-}$ and even $\Xi^{-}$ hyperon Fermi seas can be occupied at approximately the same densities, $n>(2-3)n_{0}$. The appearance of any new fermion degree of freedom decreases the energy and, consequently, the maximum compact star mass. Such a decrease of the maximum neutron star mass has been observed in many models [249, 69, 255, 72, 73]. These problems are known as the “hyperon puzzle” and the $\Delta$ puzzle”. They have been resolved in the RMF models, in which the hadron masses are assumed to be functions of the scalar $\sigma$ mean field [195, 196, 152].
VI Strangelets, strange and hybrid stars
The idea of stable strange quark matter proposed by Witten in 1984 [298] is in some sense similar to the mentioned above idea of stable $\Delta$ resonance matter by Troitsky and Chekunaev from year 1979 [269]. A quark nugget (strangelet or nuclearite) is a baryonic state composed of quarks in a deconfined phase. There is a correspondence between this state and nontopological solitons with baryon constituents studied by Lee and collaborators and then by Sidney Coleman [90, 55, 184, 178, 177]. In Ref. [298] Witten estimated the energy per quark for a three-flavor system (strange quark matter) to be $\sim 90\%$ of the energy of a two-flavour system. This additional energy gain could be enough to compensate the energy loss due to the heavier strange quark mass leading, thereby, to a stable quark system. The strange quark matter can be stable or metastable for a wide range of strong interaction parameters. If the strange quark matter is more stable than the nucleus of Fe and the baryon number is $A\ll 10^{56}-10^{57}$ one speaks about strangelets and quark nuclearites, and if the baryon number is large, $A\sim 10^{56}-10^{57}$, about strange stars. In the latter case, the gravity essentially contributes to binding. The surface layer of these objects has typically a microscopic length $\sim 10$ fm. Provided the gravity is switched off some models of quark matter do not allow for the absolutely stable strange quark matter but allow for a metastable state. Other models use such parameters that the strange and non-strange deconfined matter is unstable. In these both cases, the quark matter, nevertheless, may exist in interiors of compact stars, named in this case hybrid stars. All such stars have macroscopic hadron shells. Sometimes some authors conflate mentioned terms calling hybrid stars as strange stars or quark stars. From our point of view, the most important is that strangelets, quark nuclearites and strange stars could be stable already without assistance of gravity. The difference is only that in the case of strange stars the baryon number typically is $A\sim 10^{56}-10^{57}$, and then the gravity contributes essentially to binding. For object with a smaller $A$ the gravity is less important. However, the important difference exists between strange stars and hybrid stars, since the latter contain hadronic shells of macroscopic size. Also we should note that in all quark nuggets with $A\gg 1/e^{3}$ up to a strange star size, the net electric charge is screened mainly by electrons and muons similarly to the case of the pion condensate nuggets and nuclei-stars, cf. results of Ref. [6] and [212, 209].
For strange matter to form, there must initially be a sufficient number of strange quarks. Since the probability of multiple weak processes occurring simultaneously is enormously suppressed, the large amount of strangeness required to satisfy conditions of Witten’s hypothesis can only be obtained under very specific conditions. As far as we know, these conditions can only be met in the Early Universe and/or in the cores of massive stellar objects. Discussions of hypothetical strangelets and hybrid and strange stars began with the works [128, 127, 34, 298, 80, 115] and remain a hot topic till these days, see, e.g., Refs. [190, 296, 81, 189, 67] and references therein.
VI.1 Quark models
The most commonly employed model of quark matter is the MIT bag model [53]. By balancing the vacuum pressure outside the bag with the pressure of quarks inside it, one mimics phenomena of quark confinement and asymptotic freedom [80]. The quark matter is stable, metastable or unstable in these models in dependence mainly on the assumed value of the bag constant $B$ and the strange quark mass $m_{s}$. A large value of $B$ excludes the existence of hybrid stars. With a decrease of $B$, first the most massive neutron stars acquire a quark core, becoming hybrid stars. For a small $B$ and a smaller value of $m_{s}$, strange quark matter becomes absolutely stable, i.e., more stable than 56Fe.
In addition to the MIT-bag based approaches, the Nambu–Jona-Lasinio (NJL) model is frequently employed, see Ref. [148]. However the latter does not describe quark confinement, and its preferred sets of parameters generally do not allow for stable strange quark matter [44]. Competition of inhomogeneous chiral phases and superconducting quark phases in the NJL model was studied in [170].
The quark objects are anticipated to be color superconductors. The attraction comes either from the one-gluon exchange, or from a nonperturbative four-point interaction motivated by instantons, or from nonperturbative gluon propagators. Various quark-paring patterns were investigated: the color-flavor-locked (CFL) phase, the two-flavor color superconductivity (2SC) phase, the color spin-locking (CSL) phase, etc., see Refs. [12, 81]. The 2SC phase allows for unpaired quarks of one color. The CFL phase deals with very large values of the baryon chemical potential. In this case the color superconductivity is complete in the sense that the diquark condensation produces a gap for quarks of all three colors and flavors. The value of the gap for CFL and 2SC phases can be of the order of 100 MeV. Since the CFL phase is expected to occur at ultra-high-density, its presence in compact stars is more questionable than presence of the 2SC phase. Owing to strong interactions, a fluctuation region near the critical point of the color superconductivity is expected to be broad and it is not excluded that some effects of the color superconducting fluctuations can be manifested in heavy-ion collisions [286].
Depending on the value of the diquark coupling a region of the phase diagram was found where the inhomogeneous chiral and the 2SC condensates may coexist, see Ref. [45] and references therein. Let us add here that at large densities QCD may exhibit a moat regime in the scalar-pseudoscalar sector [226] similar to the occurrence of the minimum in the effective pion gap, see Sect. IV. Also, one uses the quark models with quark masses varying with the baryon number density, see Ref. [299]. The RMF approach proposed in Ref. [138] gives a simple way to model confinement of quarks via introducing the particular density dependence of quark masses divergent for low baryon densities and a density-dependent screening effect. The screening is described analogously to an excluded volume effect for a hadronic matter model with density-dependent coupling constants. Effects of various pasta phases occurring in the quark-hadron phase transition [281] were then investigated in Ref. [197] for a set of RMF equations of state for both the hadron matter and the quark one. Parameters used in these works allow for hybrid stars but exclude stable quark star configurations.
Some constraints can be gained from the lattice QCD calculations. The standard bag and NJL models cannot reproduce the lattice data without significant modifications. However, the lattice results at zero and small finite chemical potential are appropriately described within phenomenological quasiparticle models employing massive quarks and gluons with masses dependent on temperature [234, 106, 222, 107]. These models were also used to extrapolate the lattice equation of state to the case of cold baryon matter. This has been done in Ref. [223] and then in Refs. [224, 130, 256] and in other works.
The quasiparticle model [130] for the equation of state of the quark-gluon matter at finite temperatures and baryon chemical potential, which is tuned to the lattice QCD data [86, 58], was used in Ref. [129] to study the possibility of a hadron-to-quark transition in the neutron star matter. For the hadronic equation of state the RMF model [154] was used. For comparison two ”light-bag” and ”heavy-bag” models were also considered. The “light-bag” models with conventional current quark masses ($m_{u}=5,m_{d}=7,m_{s}=150$ MeV) and vanishing gluon mass and values of the bag constant $B^{1/4}=155$ MeV and $B^{1/4}=200$ MeV were used. These values of $B^{1/4}$ MeV are typical for models applied to the treatment of hybrid and strange stars [9, 118]. It proved to be that the “light-bag” models cannot describe the equation of state obtained in the lattice QCD calculations. Also it was considered a “heavy-bag” model with parameters $m_{u}=m_{d}=330$ MeV, $m_{s}=450$ MeV, $m_{g}=600$ MeV, and $B^{1/4}=183$ MeV. This model can reasonably reproduce the lattice results at temperatures not too close to the critical temperature. Also a more specific quark quasiparticle model [130] was introduced, which fits well the lattice QCD calculations above the critical temperature. The general result of Ref. [129] is that the families of ordinary neutron stars and pure quark stars are well separated. The hadron-quark interface can be constructed only between the light-bag model and unrealistically stiff Walecka RMF model [295]. For the quark matter model fitted to the lattice data the hybrid star cannot be constructed. The application of these models to the quark stars is illustrated in Fig. 10. The light-bag model produces neutron stars with masses up-to 1.5–2.1 $M_{\odot}$ and radii $9<R<12$ km. For the heavy-bag model, neutron star becomes lighter and more compact, and the model tuned to the lattice QCD supports only stars with very low masses ($M\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}1M_{\odot}$), high densities ($n\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10n_{0}$), and small radii ($R\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}6$ km). However the situation remains rather uncertain. Some studies provide arguments for stability or metastability of two-flavor quark matter, while others argue against the stability of quark matter, e.g. see [123, 19] and references therein.
Figure 10: Strange star mass-radius relation for the light-bag and heavy-bag models and the quasiparticle model [130] tuned to the lattice QCD data. For details see text and Ref.  [129]. The bands correspond to variations of pairing gap $\Delta$ between 0 ad 100 MeV for the CFL type of quark superconductivity.
We also should mention the possibility of a quarkyonic phase [202], which may occur between hadron and quark phases. Thus the temperature–baryon density, TnT-n, phase diagram of the nuclear matter proves to be very rich. With increasing density at a moderate temperature there may appear region, in which we may deal with the stable or metastable $\pi$ or $\sigma$ condensate matter or the strange quark matter.